The 2017 balance sheet of Kerbers Tennis Shop, Inc., showed long-term debt of $6 million, and the 2018 balance sheet showed long-term debt of $6.4 million. The 2018 income statement showed an interest expense of $160,000. What was the firms cash flow to creditors during 2018? (A negative answer should be indicated by a minus sign. Enter your answer in dollars, not millions of dollars, e.g., ,234,567.) Cash flow to creditors
Answer #1

Firm's Cash flow to Creditors = Interest expense - New debt = 160,000 -( 6,400,000 - 6,000,000) = 160,000 - 400,000 = -240,000

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well